# Phenomenology of Non-Abelian Gauge and Goldstone Bosons in a U(2) Flavor Model

### v1

@@ Abstract
We investigate the phenomenological implications of the bosons associated with the $SU(2)_{F}$ subgroup in a simple and realistic $U(2)_{F}$ flavor model. While the PNGB of the $U(1)_{F}$ factor behaves as a standard QCD axion (an axiflavon) with suppressed flavor-violating couplings, the three degrees of freedom from $SU(2)_{F}$ have not been studied before. This work focuses on these states, considering both the case where $SU(2)_{F}$ is a global symmetry, yielding pseudo-Nambu-Goldstone bosons (PNGBs), and the case where it is a gauge symmetry with a potentially small coupling, yielding a triplet of (possibly) light gauge bosons. In both scenarios, these new bosons naturally feature unsuppressed flavor-violating couplings to Standard Model fermions in the mass basis. We derive the resulting predictions for flavor-changing neutral currents and lepton flavor violation, including exotic decays of mesons and leptons. Our analysis shows that processes like $K\to\pi X$ and $\mu\to eX$ place the most stringent constraints, probing the flavor symmetry breaking scale $v_{\phi}$ up to $10^{11}-10^{12}$ GeV for light bosons, while heavier states are tested in $B$ and $\tau$ decays, as well as by $K-\bar{K}$ mixing and $\mu\to e\gamma$. We demonstrate that low-energy flavor experiments provide a powerful probe of this framework, capable of testing ultra-high symmetry breaking scales that surpass the limits set by astrophysical observations.
@@ 1 Introduction
The “strong CP problem” is the fact that, so far, no CP-violating effects stemming from the operator $\mathcal{L}\supset\bar{\theta}\frac{g_{s}}{32\pi^{2}}G\widetilde{G}$ (where $G$ and $\widetilde{G}$ are the QCD field strength and its dual, and $g_{s}$ is the strong coupling) have been observed, implying that the QCD effective angle $\bar{\theta}$ must have a tiny value, $\bar{\theta}<10^{-10}$ — see [4] for an introduction to the subject. Interestingly, one can observe a twofold connection between the strong CP problem and the origin of fermion masses, that is, the flavor puzzle: (i) a contribution to $\bar{\theta}$ arises from quark masses (see e.g. [5] for a review); (ii) the Peccei-Quinn (PQ) symmetry that can provide a dynamical solution to the problem [6, 7, 8] is a subgroup of the large SM global flavor symmetry group $U(3)^{5}$ [9]. Furthermore, the interesting possibility that another, flavor non-universal, subgroup of $U(3)^{5}$ could act as a PQ symmetry, leading to a flavor-violating axion, was suggested by Wilczek long ago [10].
@@
The simplicity and attractive features of flavor models based on a FN $U(1)$ symmetry explain why they are the object of persistent interest in the literature — see e.g. [18, 19, 20, 21, 22, 23, 24] for recent examples. Nevertheless, this class of models also has some unsatisfactory aspects. In particular: (i) the large number of free parameters — especially the $\mathcal{O}(1)$} coefficients that appear in each entry of the Yukawa matrices — limits its predictive power; (ii) typical UV completions require the introduction of a great number of fields, as systematically discussed in [25, 26]; (iii) an Abelian symmetry is unable to provide a rationale for the existence of three fermion generations.
All of these drawbacks can be addressed, or at least ameliorated, within models with a non-Abelian flavor symmetry. The simplest example — and perhaps the best motivated by the observed pattern of fermion masses — is provided by $U(2)$. Third-generation fermions (top, bottom and tau) are significantly heavier than the corresponding first two generations of quarks and leptons. Hence, the latter can be considered in first approximation massless, as they would be if belonging to doublets of an unbroken symmetry, under which the third generations are singlets. As a result, the non-Abelian $U(2)$ flavor symmetry naturally emerges as a promising candidate to describe the flavor structure observed in the data [27, 28, 29, 30, 31, 32, 33]. A simple and realistic flavour model based on $U(2)_{F}=SU(2)_{F}\times U(1)_{F}$ has been proposed in [32], with charge assignments compatible with the unification of quarks and leptons within $SU(5)$. In this framework, two flavons are introduced, an $SU(2)_{F}$ doublet and a singlet, hence the $U(2)_{F}$ symmetry breaking is controlled by two small parameters related to the VEVs of such fields. The resulting operators provide an excellent fit to all SM flavor observables, including neutrino masses and mixing angles [32, 34, 35, 36].
@@ 2 Set-up of the model
Table 1: Transformation properties of the fields under $U(2)_{F}=SU(2)_{F}\times U(1)_{F}$. $Q$ and $L$ are respectively left-handed (LH) quark and lepton doublets. $U$, $D$, $E$ are respectively right-handed (RH) up-type quark, down-type quark and lepton singlets. The subscripts denote flavor indices and, in particular, $a=1,2$ is the index of the fundamental representation of $SU(2)_{F}$. $H$ denotes the SM Higgs field, while $\phi_{a}$ and $\chi$ are the flavon fields.
In this section, we review the basic framework of the $U(2)_{F}$ flavor model [32] and introduce the Nambu-Goldstone bosons associated with the spontaneous breaking of the flavor symmetry. The transformation properties of the SM fermions under $U(2)_{F}=SU(2)_{F}\times U(1)_{F}$ are displayed in Table 1. The first two generations of fermions transform as doublets under $SU(2)_{F}$, while the third generation is an $SU(2)_{F}$ singlet. Hence, we adopt the notation $Q_{a}=(Q_{1}\ Q_{2})^{T}$ to represent the first and second generation LH quarks, and similarly for the LH leptons. The Higgs field is a singlet under $SU(2)_{F}$ and is $U(1)_{F}$-neutral.
@@
where $\lambda_{ij}^{f}$ are dimensionless coefficients (stemming from products of the fundamental couplings of the UV-complete theory) that are assumed to be $\mathcal{O}(1)$}, such that, after inserting the flavon VEVs, the Yukawa hierarchies emerge solely from powers of the small parameters $\varepsilon_{\phi}$ and $\varepsilon_{\chi}$ defined in Eq. (1). In terms of such parameters, the resulting Yukawa matrices are
@@ 3 $SU(2)_{F}$ gauge bosons
where $g_{F}$ is the $SU(2)_{F}$ gauge coupling.
@@ 4 $SU(2)_{F}$ pseudo Nambu-Goldstone bosons
where $\sigma_{3}$ denotes the third Pauli matrix. Substituting Eq. (10) into this expression, mass terms for the $SU(2)_{F}$ PNGBs arise such that $m_{\pi^{\prime}_{1}}=m_{\pi^{\prime}_{2}}=m_{\pi^{\prime}}$, $m_{\pi^{\prime}_{3}}=m_{\pi^{\prime}}/\sin\theta$ where we have taken into account the mixing ($\sin\theta\equiv v_{\chi}/\sqrt{v_{\phi}^{2}+v_{\chi}^{2}}$) with $U(1)_{F}$ PNGB as in Eq. (11). We can also assume this breaking term to arise from Planck-suppressed operators of the form $(\phi^{\dagger}\sigma^{3}\phi)(\phi^{\dagger}\phi)^{n}$ with $n>1$, such that $m_{\pi^{\prime}}^{2}\sim(v_{\phi})^{2n}/M_{\text{Pl}}^{2(n-1)}$, where $M_{\text{Pl}}\simeq 1.22\times 10^{19}\penalty 10000\ \mathrm{GeV}$ denotes the Planck scale. This assumption is consistent with the commonly accepted notion that quantum gravity effects break any global symmetry — cf. [45] for a recent review. In either case, we treat $m_{\pi^{\prime}}$ as a free parameter throughout the rest of the paper.
@@
$Q_{a}\to e^{i(\widetilde{\pi}_{1}\sigma_{1}+\widetilde{\pi}_{2}\sigma_{2}+\widetilde{\pi}^{\prime}_{3}(x)\sigma_{3})_{ab}/v_{\phi}}e^{-i\widetilde{a}/v_{\chi}}Q_{b}\quad(a,b=1,2),$ (29)
@@ 5 Phenomenological implications of the new flavored bosons
In both scenarios, flavor-violating couplings of the $U(2)_{F}$ bosons with SM fermions are naturally induced by the structure of the symmetry. In particular, $SU(2)_{F}\subset U(2)_{F}$ leads to unsuppressed flavor violation involving the first two fermion generations. In addition, flavor-violating processes have a mild dependence on the $\mathcal{O}(1)$} coefficients of the model and, in the case of those mediated or involving the axiflavon, the $U(1)_{F}$ charges of the SM fermions.
@@ 5.1 Decays into light $SU(2)_{F}$ bosons
where $L$ represents the detector size (e.g. $L\approx 1\,\mathrm{m}$ typical LFV experiments [56, 15], $L\approx 100\,\mathrm{m}$ for kaon experiments such as NA62 [57], etc.), and $l_{X}$ is the mean decay length of $X$ in the laboratory frame that, in natural units, is given by
@@
In the following, we focus on the leptonic decays $X\to\ell\ell^{(\prime)}$. $\ell,\ell^{\prime}=e,\mu$, which are the main allowed channels in the regime $m_{X}\lesssim 1\,\mathrm{GeV}$ and the cleanest otherwise. For a given value of its mass $m_{X}$, the average decay length of $X$ depends on its couplings to fermions. If the latter are sufficiently large, $X$ will decay promptly within the detector, producing visible dilepton signatures. In contrast, for small couplings, $X$ may be long-lived and escape the detector before decaying, resulting in a missing energy signal.
@@
$K\to\pi\,X_{\text{inv}}\,,\quad B\to K(\pi)\,X_{\text{inv}}\,,\quad\mu\to\,X_{\text{inv}}\,,\quad\tau\to\ell\,X_{\text{inv}}\,$, (42)
@@
$\left[1-\exp\left(-\frac{L}{l_{X}}\right)\right]\,\mathrm{BR}(P\to P^{\prime}\,X)\,\mathrm{BR}(X\to\ell\ell^{(\prime)})\penalty 10000\ <\penalty 10000\ \mathrm{BR}(P\to P^{\prime}\,\ell\ell^{(\prime)})^{\text{exp}}\,$, (45)
that enable us to confront our model’s predictions with existing experimental bounds on processes with both visible and invisible final states — bounds that are respectively denoted as $\mathrm{BR}(P\to P^{\prime}+\text{inv})^{\text{exp}}$ and $\mathrm{BR}(P\to P^{\prime}\,\ell\ell^{(\prime)})^{\text{exp}}$. The experimental limits we use in our phenomenological study are presented in Appendix B.
@@ 5.2 Flavor violation mediated by heavy $SU(2)_{F}$ bosons
where the loop functions $g$ and $\widetilde{g}$ are listed in the Appendix A. Despite the $\sim m_{\tau}/m_{\mu}$ enhancement, the second terms in the above expressions give contributions of the same order as (in fact smaller than) the first ones, because flavor-changing interactions involving the third generation are parametrically smaller due to the $SU(2)_{F}$ coupling structure. Similarly, we can derive expressions for $\tau\to\ell\gamma$ ($\ell=\mu,e$) adapting the formulae in [55] and taking into account that, in this case, $\tau$ loops are always subdominant because of the large suppression of the flavor-conserving coupling to the third generation, $\hat{C}^{e\,i}_{A\,\tau\tau}$. Hence, the resulting constraints are not competitive with other limits, in particular from $\mu\to e\gamma$:
@@ 5.3 Numerical results and discussion
For comparison, Figures 1 and 2 also show astrophysical bounds (cyan dashed lines) estimated as explained in Appendix C. The most relevant constraint to the couplings of the light fields with quarks is derived from the SN1987A explosion data, while the couplings to electrons are constrained by white dwarfs (WD) and red giants (RG) cooling and, again, supernova explosion — the latter bound is labeled as SN1987A(ee) in the right plots. WD and RG bounds are the most stringent ($v_{\phi}\gtrsim{\mathcal{O}(10^{9})}\penalty 10000\ \text{GeV}$) but they do not apply to particles with $m_{X}\gtrsim 10-100$ keV, that is, too much above the average temperature of the star plasma to be efficiently produced. Hence, they are outside the parameter range displayed in the figures. The supernova bounds exhibit the typical Boltzmann suppression for $m_{X}\gtrsim 100\,\text{MeV}$ while, for simplicity, we do not calculate the loss of sensitivity in the large coupling regime when the light particles remain trapped in the star and do not contribute to its cooling. In any case, these estimates demonstrate that these astrophysical limits on light $SU(2)_{F}$ bosons are subdominant compared to the bounds from laboratory experiments. More important are the limits one can obtain from axiflavon emission from the star plasma, which are indicated as gray dashed lines in the figures. Recall that, in both scenarios we consider, this field is the PNGB of the global factor $U(1)_{F}\subset U(2)_{F}$ and behaves like a QCD axion acquiring a tiny mass — see Eq. (17) — from the QCD anomaly, hence it is always light enough to be produced in the star plasma. The most relevant resulting constraint (from SN1987A cooling) translates to a limit $v_{\phi}\gtrsim 3\times 10^{9}\penalty 10000\ \text{GeV}$ that, however strong, does not prevent the observation of light $SU(2)_{F}$ bosons in kaon or muon decays. Furthermore, in the large coupling/small $v_{\phi}$ region also the axiflavon enters the trapping regime, which makes the other processes discussed in this section relevant tests of such a portion of the parameter space.

### v2

@@ Abstract
We investigate the phenomenological implications of the bosons associated with the $SU(2)_{F}$ subgroup in a simple and realistic $U(2)_{F}$ flavor model. While the Nambu-Goldstone boson of the $U(1)_{F}$ factor behaves as a standard QCD axion (an axiflavon) with suppressed flavor-violating couplings, the three degrees of freedom from $SU(2)_{F}$ have not been studied before. This work focuses on these states, considering both the case where $SU(2)_{F}$ is a global symmetry, yielding pseudo-Nambu-Goldstone bosons (PNGBs), and the case where it is a gauge symmetry with a potentially small coupling, yielding a triplet of (possibly) light gauge bosons. In both scenarios, these new bosons naturally feature unsuppressed flavor-violating couplings to Standard Model fermions in the mass basis. We derive the resulting predictions for flavor-changing neutral currents and lepton flavor violation, including exotic decays of mesons and leptons. Our analysis shows that processes like $K\to\pi X$ and $\mu\to eX$ place the most stringent constraints, probing the flavor symmetry breaking scale $v_{\phi}$ up to $10^{11}-10^{12}$ GeV for light bosons, while heavier states are tested in $B$ and $\tau$ decays, as well as by $K-\bar{K}$ mixing and $\mu\to e\gamma$. We demonstrate that low-energy flavor experiments provide a powerful probe of this framework, capable of testing ultra-high symmetry breaking scales that surpass the limits set by astrophysical observations.
@@ 1 Introduction
The “strong CP problem” is the fact that, so far, no CP-violating effects stemming from the operator $\mathcal{L}\supset\bar{\theta}\frac{g_{s}}{32\pi^{2}}G\widetilde{G}$ (where $G$ and $\widetilde{G}$ are the QCD field strength and its dual, and $g_{s}$ is the strong coupling) have been observed, implying that the QCD effective angle $\bar{\theta}$ must have a tiny value, $\bar{\theta}<10^{-10}$ — see [4] for an introduction to the subject. Interestingly, one can observe a twofold connection between the strong CP problem and the origin of fermion masses, that is, the flavor puzzle: (i) a contribution to $\bar{\theta}$ arises from quark masses (see e.g. [5] for a review); (ii) the Peccei-Quinn (PQ) symmetry that can provide a dynamical solution to the problem [6, 7, 8] is a subgroup of the large SM global flavor symmetry group $U(3)^{5}$ [9]. Furthermore, the interesting possibility that another, flavor non-universal, subgroup of $U(3)^{5}$ could act as a PQ symmetry, leading to a flavor-violating axion, was suggested by Wilczek long ago [10] — see also [11, 12, 13, 14, 15, 16, 17, 18, 19] for related early discussions on Nambu-Goldstone bosons from flavor symmetries, including non-Abelian examples.
@@
The simplicity and attractive features of flavor models based on a FN $U(1)$ symmetry explain why they are the object of persistent interest in the literature — see e.g. [27, 28, 29, 30, 31, 32, 33, 34] for recent examples. Nevertheless, this class of models also has some unsatisfactory aspects. In particular: (i) the large number of free parameters — especially the $\mathcal{O}(1)$} coefficients that appear in each entry of the Yukawa matrices — limits its predictive power; (ii) typical UV completions require the introduction of a great number of fields, as systematically discussed in [35, 36]; (iii) an Abelian symmetry is unable to provide a rationale for the existence of three fermion generations.
All of these drawbacks can be addressed, or at least ameliorated, within models with a non-Abelian flavor symmetry. The simplest example — and perhaps the best motivated by the observed pattern of fermion masses — is provided by $U(2)$. Third-generation fermions (top, bottom and tau) are significantly heavier than the corresponding first two generations of quarks and leptons. Hence, the latter can be considered in first approximation massless, as they would be if belonging to doublets of an unbroken symmetry, under which the third generations are singlets. As a result, the non-Abelian $U(2)$ flavor symmetry naturally emerges as a promising candidate to describe the flavor structure observed in the data [37, 38, 39, 40, 41, 42, 43, 44, 45]. A simple and realistic flavour model based on $U(2)_{F}\simeq SU(2)_{F}\times U(1)_{F}$ has been proposed in [42], with charge assignments compatible with the unification of quarks and leptons within $SU(5)$. In this framework, two flavons are introduced, an $SU(2)_{F}$ doublet and a singlet, hence the $U(2)_{F}$ symmetry breaking is controlled by two small parameters related to the VEVs of such fields. The resulting operators provide an excellent fit to all SM flavor observables, including neutrino masses and mixing angles [42, 46, 47, 48].
@@ 2 Set-up of the model
Table 1: Transformation properties of the fields under $U(2)_{F}\simeq SU(2)_{F}\times U(1)_{F}$. $Q$ and $L$ are respectively left-handed (LH) quark and lepton doublets. $U$, $D$, $E$ are respectively right-handed (RH) up-type quark, down-type quark and lepton singlets. The subscripts denote flavor indices and, in particular, $a=1,2$ is the index of the fundamental representation of $SU(2)_{F}$. $H$ denotes the SM Higgs field, while $\phi_{a}$ and $\chi$ are the flavon fields.
In this section, we review the basic framework of the $U(2)_{F}$ flavor model [42] and introduce the Nambu-Goldstone bosons associated with the spontaneous breaking of the flavor symmetry. The transformation properties of the SM fermions under $U(2)_{F}\simeq SU(2)_{F}\times U(1)_{F}$ are displayed in Table 1. The first two generations of fermions transform as doublets under $SU(2)_{F}$, while the third generation is an $SU(2)_{F}$ singlet. Hence, we adopt the notation $Q_{a}=(Q_{1}\ Q_{2})^{T}$ to represent the first and second generation LH quarks, and similarly for the LH leptons. The Higgs field is a singlet under $SU(2)_{F}$ and is $U(1)_{F}$-neutral.
@@
where $\lambda_{ij}^{f}$ are dimensionless coefficients (stemming from products of the fundamental couplings of the UV-complete theory) that are assumed to be $\mathcal{O}(1)$}, such that, after inserting the flavon VEVs, the Yukawa hierarchies emerge solely from powers of the small parameters $\varepsilon_{\phi}$ and $\varepsilon_{\chi}$ defined in Eq. (1). In particular, notice that the above operators, hence the Yukawa interactions, do not depend on the absolute scale of the flavon VEVs nor on the cutoff scale $\Lambda$. Thus, these scales cannot be determined in terms of measured fermion masses and mixing but can only be constrained by flavour processes, as discussed in the following sections. In terms of the expansion parameters and the $\mathcal{O}(1)$} coefficients, the Yukawa matrices result:
@@ 3 $SU(2)_{F}$ gauge bosons
where $g_{F}$ is the $SU(2)_{F}$ gauge coupling. Hence, we see that $W^{\prime}$ comprises three degenerate electromagnetically neutral states.
@@ 4 $SU(2)_{F}$ pseudo Nambu-Goldstone bosons
where $\sigma_{3}$ denotes the third Pauli matrix. Substituting Eq. (10) into this expression, mass terms for the $SU(2)_{F}$ PNGBs arise such that $m_{\pi^{\prime}_{1}}=m_{\pi^{\prime}_{2}}=m_{\pi^{\prime}}$, $m_{\pi^{\prime}_{3}}=m_{\pi^{\prime}}/\sin\theta$ where we have taken into account the mixing ($\sin\theta\equiv v_{\chi}/\sqrt{v_{\phi}^{2}+v_{\chi}^{2}}$) with $U(1)_{F}$ PNGB as in Eq. (11). We can also assume this breaking term to arise from Planck-suppressed operators of the form $(\phi^{\dagger}\sigma_{3}\phi)(\phi^{\dagger}\phi)^{n}$ with $n>1$, such that $m_{\pi^{\prime}}^{2}\sim(v_{\phi})^{2n}/M_{\text{Pl}}^{2(n-1)}$, where $M_{\text{Pl}}\simeq 1.22\times 10^{19}\penalty 10000\ \mathrm{GeV}$ denotes the Planck scale. This assumption is consistent with the commonly accepted notion that quantum gravity effects break any global symmetry — cf. [57] for a recent review. In either case, we treat $m_{\pi^{\prime}}$ as a free parameter throughout the rest of the paper.
@@
$Q_{a}\to e^{i(\widetilde{\pi}_{1}\sigma_{1}+\widetilde{\pi}_{2}\sigma_{2}+\widetilde{\pi}^{\prime}_{3}\sigma_{3})_{ab}/v_{\phi}}e^{-i\widetilde{a}/v_{\chi}}Q_{b}\quad(a,b=1,2),$ (29)
@@ 5 Phenomenological implications of the new flavored bosons
In both scenarios, flavor-violating couplings of the $U(2)_{F}$ bosons with SM fermions are naturally induced by the structure of the symmetry. In particular, $SU(2)_{F}\subset U(2)_{F}$ leads to unsuppressed flavor violation involving the first two fermion generations. In addition, flavor-violating processes have a mild dependence on the $\mathcal{O}(1)$} coefficients of the model and, in the case of those mediated or involving the axiflavon, the $U(1)_{F}$ charges of the SM fermions. [^6] In order to evaluate a given process, the couplings appearing in the Lagrangians in the previous sections should be evolved from the symmetry-breaking scale down to the energy scale characteristic of the process of interest. The renormalization group equations for the couplings of axions and ALPs are, for instance, given and their numerical impact is studied in [58, 59, 60]. However, for the flavor-violating interactions that we are interested in, running effects are suppressed by small Yukawa couplings and/or CKM angles. In agreement with the numerical estimates in [59], we find that such corrections are thus always below the 10% level, hence we neglect them throughout our analysis.
@@ 5.1 Decays into light $SU(2)_{F}$ bosons
where $L$ represents the detector size (e.g. $L\approx 1\,\mathrm{m}$ for typical LFV experiments [71, 72], $L\approx 100\,\mathrm{m}$ for kaon experiments such as NA62 [73], etc.), and $l_{X}$ is the mean decay length of $X$ in the laboratory frame that, in natural units, is given by
@@
In the following, we focus on the leptonic decays $X\to\ell\ell^{(\prime)}$. $\ell,\ell^{\prime}=e,\mu$, which are the main allowed channels in the regime $m_{X}\lesssim 1\,\mathrm{GeV}$ and the cleanest otherwise. For a given value of its mass $m_{X}$, the average decay length of $X$ depends on its couplings to fermions. If the latter are sufficiently large, $X$ will decay promptly within the detector, producing visible dilepton signatures. In contrast, for small couplings, $X$ may be long-lived and escape the detector before decaying, resulting in a missing energy signal. [^7] The viability and potential impact of the intermediate regime — corresponding to displaced-vertex decays — strongly depend on the specific process and detector. See Appendix B for more detailed comments.
@@
$K\to\pi\,X_{\text{inv}}\,,\quad B\to K(\pi)\,X_{\text{inv}}\,,\quad\mu\to e\,X_{\text{inv}}\,,\quad\tau\to\ell\,X_{\text{inv}}\,$, (42)
@@
$\left[1-\exp\left(-\frac{L_{\text{prompt}}}{l_{X}}\right)\right]\,\mathrm{BR}(P\to P^{\prime}\,X)\,\mathrm{BR}(X\to\ell\ell^{(\prime)})\penalty 10000\ <\penalty 10000\ \mathrm{BR}(P\to P^{\prime}\,\ell\ell^{(\prime)})^{\text{exp}}\,$, (45)
where $L_{\text{prompt}}$ represents the decay length under which the decay appears prompt, which depends on the spatial resolution of a given detector. For simplicity, we take $L_{\text{prompt}}=1\,\mathrm{mm}$ throughout our analysis with the exception of kaon processes, for which it is sufficient that $X$ decays within the detector to give rise to a visible signature. The above formulae enable us to confront our model’s predictions with existing experimental bounds on processes with both visible and invisible final states — bounds that are respectively denoted as $\mathrm{BR}(P\to P^{\prime}+\text{inv})^{\text{exp}}$ and $\mathrm{BR}(P\to P^{\prime}\,\ell\ell^{(\prime)})^{\text{exp}}$. The experimental limits we use in our phenomenological study are presented in Appendix B.
@@ 5.2 Flavor violation mediated by heavy $SU(2)_{F}$ bosons
where the loop functions $g$ and $\widetilde{g}$ are listed in the Appendix A. Despite the $\sim(m_{\tau}/m_{\mu})^{2}$ enhancement, the second terms in the above expressions give contributions of the same order as (in fact smaller than) the first ones, because flavor-changing interactions involving the third generation are parametrically smaller due to the $SU(2)_{F}$ coupling structure. Similarly, we can derive expressions for $\tau\to\ell\gamma$ ($\ell=\mu,e$) adapting the formulae in [70] and taking into account that, in this case, $\tau$ loops are always subdominant because of the large suppression of the flavor-conserving coupling to the third generation, $\hat{C}^{e\,i}_{A\,\tau\tau}$. Hence, the resulting constraints are not competitive with other limits, in particular from $\mu\to e\gamma$:
@@ 5.3 Numerical results and discussion
For comparison, Figures 1 and 2 also show astrophysical bounds (cyan dashed lines) estimated as explained in Appendix C. The most relevant constraint to the couplings of the light fields with quarks is derived from the SN1987A explosion data, while the couplings to electrons are constrained by white dwarfs (WD) and red giants (RG) cooling and, again, supernova explosion — the latter bound is labeled as SN1987A(ee) in the right plots. WD and RG bounds are the most stringent ($v_{\phi}\gtrsim{\mathcal{O}(10^{9})}\penalty 10000\ \text{GeV}$) but they do not apply to particles with $m_{X}\gtrsim 10-100$ keV, that is, too much above the average temperature of the star plasma to be efficiently produced. Hence, they are outside the parameter range displayed in the figures. The supernova bounds exhibit the typical Boltzmann suppression for $m_{X}\gtrsim 100\,\text{MeV}$ while, for simplicity, we do not calculate the loss of sensitivity in the large coupling regime when the light particles remain trapped in the star and do not contribute to its cooling. In any case, these estimates demonstrate that these astrophysical limits on light $SU(2)_{F}$ bosons are subdominant compared to the bounds from laboratory experiments. More important are the limits one can obtain from axiflavon emission from the star plasma, which are indicated as gray dashed lines in the figures. Recall that, in both scenarios we consider, this field is the PNGB of the global factor $U(1)_{F}\subset U(2)_{F}$ and behaves like a QCD axion acquiring a tiny mass — see Eq. (17) — from the QCD anomaly, hence it is always light enough to be produced in the star plasma — see Section 2.2 and Appendix D. The most relevant resulting constraint (from SN1987A cooling) translates to a limit $v_{\phi}\gtrsim 3\times 10^{9}\penalty 10000\ \text{GeV}$ that, however strong, does not prevent the observation of light $SU(2)_{F}$ bosons in kaon or muon decays, as illustrated by Figures 1 and 2. Furthermore, in the large coupling/small $v_{\phi}$ region also the axiflavon enters the trapping regime, which makes the other processes discussed in this section relevant tests of such a portion of the parameter space.